"One salt solution is 5% salt and a second is 20% salt. How many ounces of each should be mixed to make one quart of 14% salt solution?"

Answer by stanbon(75887)

One salt solution is 5% salt and a second is 20% salt. How many ounces of each should be mixed to make one quart of 14% salt solution
----------------------
Equation:
salt + salt = salt
0.05x + 0.20(1-x) = 0.14*1
----
Multiply thru by 100 to get:
5x + 20(1-x) = 14
5x + 20 - 20x = 14
-15x = -6
x = 2/5 = 0.4 quarts (amt. of 5% solution needed)
----------
1-x = 0.6 quarts (amt. of 20% solution needed)
==================================
Cheers,
Stan H.
=============